A bacterium moves with a speed of 1.6 µm/s across a petri dish with radius
r = 9.1 cm.
How long does it take the bacterium to traverse the petri dish along its diameter?
1
Expert's answer
2019-02-05T11:50:52-0500

We can find the time that the bacterium takes to traverse the petri dish along its diameter from the formula:

"t = \frac{d}{v} = \frac{2r}{v},"

here,

"d = 2r"

is the diameter of the petri dish,

"r"

is the radius of the petri dish,

"v"

is the speed of the bacterium.

Then, we get:

"t = \frac{2r}{v} = \frac{2 \cdot 9.1 cm \cdot \frac{1 m}{100 cm}}{1.6 \frac{\mu m}{s} \cdot \frac{1m}{10^6 \mu m}} = 1.14 \cdot 10^5 s."

Answer:

"t = 1.14 \cdot 10^5 s"

.